A car accelerates uniformly from 12.5m/s at 10.0 degrees South of West to 105.0 m/s at 85.0 degrees South of West in 7.00 s. Determine the x and y components of the vehicles displacement. Determine the magnitude and direction of the vehicles displacement vector.


1
Expert's answer
2021-10-25T10:03:11-0400

"v_2=v_1+at,"

"a=\frac{v_2-v_1}t=13.2~\frac ms,"

"|\Delta \vec v|=|\vec v_2-\vec v_1|=102.5~\frac ms,"

"l=\Delta v t=717.5~m,"

"\beta=75°+\arcsin(\frac{12.5}{102.5}\sin 75°)=82°" south of east,

"l_x=l\cos82°=100~m,"

"l_y=l\sin 82°=710.5~m."